# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 4.2.3 Computing Power Network and AI
The NET4AI architecture requires powerful computing service capabilities because wireless networks show great interest in computing tasks related to high-performance computing, such as ultra-large-scale data processing and deep learning. The computing power network is a key enabling technology for computing services[222]. It achieves ubiquitous computing interconnection through mutual sense and collaboration between the network and computing resources[223], promotes efficient harmony between cloud, edge and terminal resources, and serves all computing tasks in the network, especially AI-related computing services.
@@
In summary, the computing power network provides stable and reliable computing resources for AI, using idle computing power in the network to complete the computing requirements of large-scale AI tasks, improving computing efficiency and reducing costs. Users do not need to build their computing infrastructure, enhancing the accessibility of AI computing and promoting the widespread application and adoption of AI technology. The computing power network integrates distributed computing resources to provide efficient, flexible, and scalable computing power to meet the real-time and environmental requirements of modern computing environments. It supports users to dynamically adjust resource scale according to demand to adapt to AI applications of different scales and complexities[227]. For example, users can request a large amount of idle computing power through the computing power network to support the high computing power requirements in new applications such as virtual reality (XR) and augmented reality (AR).
@@
With the explosive growth of data volume in the network and computing requirements becoming increasingly complex, the computing power network has become the focus of attention of many researchers to solve these complex challenges. The authors of [230] implemented a Kubernete-based prototype testing platform for the computing power network using a microservices architecture, achieving key enabling technologies for the computing power network, including computational modeling, sense, announcement, and offloading. The authors of [231] described a lightweight multi-cluster hierarchical edge resource scheduling scheme based on a cloud-native resource scheduling mechanism, successfully managing and deploying many heterogeneous edge devices within a unified framework using a lightweight cloud-native platform. The authors of [225] proposed an integrated ICT network architecture of “connectivity + computing + intelligence,” integrating network, computing, storage, and applications into a unified design. The network can sense computational resources and perform related management and control. The authors of [232] designed an on-demand compute resource scheduling model and proposed a two-stage evolutionary search scheme. Numerical results indicate that this scheme significantly improves the efficiency and stability of task offloading and compute resource scheduling in edge computing networks. Detailed power analysis of several INC use cases in [228] showed that INC becomes more energy-efficient at very low processing loads, with increased processing loads having little impact on INC power consumption. The computing power network provides support and empowerment for AI in many ways. In [233], the fog computing network offers powerful computing power support for AI model training, reasoning and other tasks. This helps shorten model training cycles, improve training efficiency, and enable AI models to iterate and optimize faster. The study presented in [234] takes deep neural networks as a typical AI application, and formulates optimization problems under the constraints of energy consumption, delay, computing and communication resources to efficiently utilize computing resources and avoid resource waste.
@@ 4.2.4 Secure Multi-Party Computation
The 6G network supports ubiquitous interconnection services. However, this will generate a large amount of user data, posing risks of privacy exposure and unauthorized use during collection, storage, and transmission. Moreover, the 6G network deeply integrates the digital and physical worlds, making social management, economic production, and daily life increasingly reliant on reliable network operations. A large amount of critical data related to social management, financial activities, and personal privacy information will be transmitted and stored through the network. The massive data flow across multiple networks and entities requires a secure, reliable, and trustworthy governance system capable of providing decentralized mechanisms and privacy protection. The 6G network will also be highly complex, heterogeneous, and dynamically changing. Extensive data exchange and collaboration between devices, service providers, and applications will be necessary. While achieving efficient cooperation, the 6G network must also ensure data security.
@@
Network attack methods are becoming increasingly sophisticated as network technology continues to evolve. Traditional security mechanisms are often inadequate to counter these emerging threats. As shown in Figure 21, secure Multi-Party Computation (SMPC) [235] is an encryption protocol that allows multiple parties to compute a function while keeping their input data private jointly. Each party obtains the correct computation result without knowing the other parties’ private inputs. SMPC has characteristics such as decentralization, input data security, and accurate computation results. It enables AI to handle complex network environments, protect user privacy and data security, support emerging application scenarios, and promote data sharing and utilization.
Figure 21: Secure Multi-Party Computation
@@
The leading supporting technologies of SMPC include Garbled Circuits [236], Oblivious Transfer [237], Secret Sharing [238], and Homomorphic Encryption [239].
@@
(1) Garbled Circuits This technique involves compiling computational logic into a circuit and encrypting each gate (such as AND, XOR, etc.). Participants can complete the computation by interacting with the encrypted information without knowing the specific logic of the circuit.
@@
(2) Oblivious Transfer Oblivious Transfer is a protocol that allows one party (e.g., the sender) to send one of many pieces of information to another (the receiver), where the receiver can choose only one piece of information. The sender does not know which piece of information the receiver has chosen.
@@
(3) Secret Sharing This method involves splitting a secret into multiple shares and distributing them to different participants. The original secret can only be reconstructed when sufficient shares are combined.
@@
(4) Homomorphic Encryption Homomorphic Encryption is a unique encryption method that allows specific computations (such as addition and multiplication) to be performed on encrypted data. When decrypted, the result of these computations corresponds to the exact computations performed on the original data.
@@
With the development of 6G networks and AI technology, SMPC provides solid guarantees for privacy protection in 6G networks based on the technologies mentioned above. It has been widely applied in privacy-preserving machine learning and other related areas. Recent research has leveraged SMPC for secure FL model aggregation. FL involves training machine learning models on numerous decentralized devices while localizing the training data. This collaborative approach allows multiple parties to train models without exposing their training data. For example, to counter reverse attacks on data, a privacy-preserving data aggregation mechanism based on secret sharing techniques in SMPC can efficiently protect FL against such attacks [240]. To address the issues of high communication overhead and poor scalability in traditional SMPC, a two-phase SMPC-based FL framework has been proposed [241]. This framework enables multiple clients to train AI models while collaboratively protecting their data privacy. To mitigate the high communication costs and poor scalability associated with aggregating global models at only some nodes, a hierarchical FL architecture using SMPC has been developed. This architecture improves communication efficiency and scalability without compromising privacy [242].
To protect the privacy of medical data and ensure it is unaffected by IoT devices, an FL scheme based on homomorphic re-encryption has been proposed [243]. In this scheme, data is collected by IoT devices, encrypted, and uploaded to fog nodes. In coordination with cloud servers, the fog nodes complete the model training. Scholars have combined homomorphic encryption with FL algorithms to protect medical data privacy from the perspective of deep learning model attacks. Using secure multi-party protocols safeguards CNN network models from malicious attacks, thus protecting real medical datasets [244]. Additionally, with encryption technology, the more complex the homomorphic algorithm used, the more training time FL consumes, indicating that the complexity of encryption technology is a significant research factor [245].
@@
In urban computing scenarios, data is often characterized by high volume, high frequency, and asynchronous communication, presenting new challenges for implementing FL. A novel hybrid FL architecture has been proposed, which combines FL with Trusted Execution Environments, SMPC and Beidou satellites. This architecture achieves secure key distribution, encryption, and decryption and provides verification mechanisms for each participant to ensure the security of local data [246]. SMPC can also be combined with blockchain technology to track and mitigate malicious behavior, achieving a secure and trustworthy collaborative edge learning framework [247]. Additionally, in privacy-preserving machine learning, SMPC can identify spam messages [248] and train text classifiers [249].
In general, SMPC enhances the security of AI models during data sharing and collaboration, ensuring data privacy. It can be combined with various AI algorithms or models, offering broad application prospects, flexibility, and scalability.
@@ 4.2.5 Semantic Communication
Semantic communication can be viewed as a key enabling technology for AI services using wireless communication services due to its higher compression and spectral efficiency, depicted in Figure 22. 6G networks will provide users with immersive, personalized, and all-encompassing services, ultimately achieving the goal of services that adapt to user desires. Under this background, networks are expected to change according to needs. With the emergence and development of new technologies, such as brain-computer interaction, brain-like computing, and semantic perception and recognition, 6G networks will possess semantic perception, recognition, analysis, understanding, and reasoning capabilities. To meet the new demands of 6G applications and services, semantic communication stands out as a paradigm that sends only meanings of semantics, addressing the latest requirements of 6G networks.
6G networks will deeply integrate multi-modal semantic perception and communication, fully utilizing standard semantic information and universal knowledge domains under different conditions, including users, devices, contexts, scenes, and environments. This integration will automatically perceive, recognize, extract, infer, and transfer the semantics and knowledge contained in transmitted information. This fundamentally solves incompatibility issues and poor interoperability across systems, protocols, networks, and human-machine interactions in traditional data-driven communication protocols, significantly improving communication efficiency, reducing information transmission and understanding delays, minimizing semantic distortion, and enhancing user quality of experience.
Figure 22: Semantic Communication
Traditional transmission methods focus on the accurate transmission of information symbols (e.g., bits), assuming that the sender and receiver are unintelligent entities, thereby overlooking the relevance and effectiveness of the transmitted contextual information. This results in the transmission of large amounts of redundant data, wasting communication resources such as transmission power and bandwidth. Moreover, acquiring, processing, and sending these vast amounts of redundant real-time data increases computational waste, communication delay, and security issues.
Semantic communication interprets information at the semantic level, transmits meanings, and restores information. Its transmitter can understand and extract semantics from information, while the receiver can receive and understand semantics to recover the data or execute the tasks. In recent years, the unique advantages of semantic communication networks in knowledge sharing and semantic understanding have gained attention in academia and have been applied in various scenarios, including tactile internet, holographic communication, XR, and human-machine symbiotic networks. Semantic communication, using semantics-oriented compression algorithms, can more effectively eliminate redundant information, significantly reducing the amount of data to be transmitted, thus improving communication efficiency and lowering bandwidth requirements. Additionally, semantic communication can understand the data context, preserving essential semantic relationships when transmitting information, thereby aiding downstream AI systems in making more intelligent and accurate decisions. It can also semantically integrate different modalities (e.g., text, images and audio), providing more comprehensive and precise information input and enhancing the collaborative capabilities of AI systems. Additionally, semantic communication offers more robust security through data anonymization, reducing the risk of data breaches.
@@
The authors of [250] provided a pioneering work in semantic information theory, establishing the mathematical framework for semantic information theory, which lays the foundation for semantic communication. The authors of [251] proposed the concept of Modulation Division Multiple Access technology, which efficiently distinguishes users based on semantic features from the information dimension, thereby improving resource utilization in multi-user semantic communication systems, which contributes to multi-user semantic communication systems. Semantic extraction and recovery, also called semantic encoding and decoding, are the core components of semantic transceivers. Four dominant advanced semantic extraction techniques [252]: DL-based semantic extraction, RL-based semantic extraction, KB-assisted semantic extraction, and native semantic extraction. Moreover, semantic communication systems (SemCom) can be categorized into text transmission, audio transmission and recognition, image transmission and recognition, and video transmission and recognition from the perspective of data modality and task classification.
@@
Text Transmission: The authors of [253] developed a DL-based text semantic communication system named DeepSC, which performs joint semantic-channel coding to achieve superior performance gains compared to existing technologies, particularly at low SNRs. Recognizing that most semantic metrics are non-differentiable, the authors of [254] introduced a reinforcement learning-based optimization paradigm for large-scale and complex text semantic transmission. This technique uses self-critical random iterative updates to train decoupled semantic transceivers, addressing the non-differentiable semantic channel optimization issue. The authors of [255] proposed a text semantic communication technique named semantic coding Reed Solomon HARQ by combining hybrid automatic repeat request (HARQ) and Reed-Solomon channel coding with semantic coding to tackle inefficiency and inflexibility with varying sentence lengths.
@@
Audio Transmission and Recognition: A DL-based audio semantic communication system named DeepSC-S was rapidly developed, as noted in [256, 257]. This system enhances transmission efficiency by transmitting only semantic information. The audio SemCom scheme employs a joint semantic encoder/decoder and channel encoder/decoder to facilitate efficient learning and voice feature extraction while mitigating wireless channel distortions. Additionally, the authors of [258] explored using FL for audio semantic communication, developing a wav2vec-based autoencoder composed of CNNs. This system effectively encodes, transmits, and decodes audio semantic information, reducing communication overhead.
@@
Image Transmission and Recognition: The authors of [259] investigated a DeepJSCC image SemCom architecture, where the encoder and decoder functions are parameterized by CNNs and jointly trained on the same dataset to minimize the mean squared error of the reconstructed images. Joint training enables DeepJSCC to avoid the cliff effect and adapt smoothly to changes in channel signal-to-noise ratio (SNR), with performance that degrades gracefully as the SNR deviates from the assumed value during training.
@@
Video Transmission and Recognition: The authors of [260] developed a video SemCom system that benefits from semantic information. This video SemCom framework utilizes nonlinear transformations and conditional coding architectures to adaptively extract the semantic features of video frames transmitted through a set of variable-length DeepJSCC encoders/decoders and wireless channels. This video SemCom technology significantly outperforms traditional wireless video coding transmission schemes.
@@
Semantic communication networks have unique advantages in knowledge sharing and semantic understanding, driving advancements in applications such as the tactile internet, holographic communication, XR, and human-machine symbiotic networks, fostering an increasingly intelligent and efficient society.

### v3

@@ 3.2.3 Computing Power Network and AI
The NET4AI architecture requires powerful computing service capabilities because wireless networks show great interest in computing tasks related to high-performance computing, such as ultra-large-scale data processing and DL. The computing power network is a key enabling technology for computing services[192]. It achieves ubiquitous computing interconnection through mutual sense and collaboration between the network and computing resources[193], promotes efficient harmony between cloud, edge and terminal resources, and serves all computing tasks in the network, especially AI-related computing services.
@@
In summary, the computing power network provides stable and reliable computing resources for AI, using idle computing power in the network to complete the computing requirements of large-scale AI tasks, improving computing efficiency and reducing costs. Users do not need to build their computing infrastructure, enhancing the accessibility of AI computing and promoting the widespread application and adoption of AI technology. The computing power network integrates distributed computing resources to provide efficient, flexible, and scalable computing power to meet the real-time and environmental requirements of modern computing environments. It supports users to dynamically adjust resource scale according to demand to adapt to AI applications of different scales and complexities[197]. For example, users can request a large amount of idle computing power through the computing power network to support the high computing power requirements in new applications such as XR and AR.
@@
With the explosive growth of data volume in the network and computing requirements becoming increasingly complex, the computing power network has become the focus of attention of many researchers to solve these complex challenges. The authors of [200] implemented a Kubernete-based prototype testing platform for the computing power network using a microservices architecture, achieving key enabling technologies for the computing power network, including computational modeling, sense, announcement, and offloading. The authors of [201] described a lightweight multi-cluster hierarchical edge resource scheduling scheme based on a cloud-native resource scheduling mechanism, successfully managing and deploying many heterogeneous edge devices within a unified framework using a lightweight cloud-native platform. The authors of [195] proposed an integrated ICT network architecture of “connectivity + computing + intelligence,” integrating network, computing, storage, and applications into a unified design. The network can sense computational resources and perform related management and control. The authors of [202] designed an on-demand compute resource scheduling model and proposed a two-stage evolutionary search scheme. Numerical results indicate that this scheme significantly improves the efficiency and stability of task offloading and compute resource scheduling in edge computing networks. Detailed power analysis of several INC use cases in [198] showed that INC becomes more energy-efficient at very low processing loads, with increased processing loads having little impact on INC power consumption. The computing power network provides support and empowerment for AI in many ways. In [203], the fog computing network offers powerful computing power support for AI model training, reasoning and other tasks. This helps shorten model training cycles, improve training efficiency, and enable AI models to iterate and optimize faster. The study presented in [204] takes DNNs as a typical AI application, and formulates optimization problems under the constraints of energy consumption, delay, computing and communication resources to efficiently utilize computing resources and avoid resource waste.
@@ 3.2.4 Secure Multi-Party Computation
Ubiquitous interconnected services will generate vast amounts of user data, which face risks of privacy leakage and illegal usage during the processes of collection, storage, and transmission. In highly complex, heterogeneous, and dynamically evolving network environments, it is imperative to ensure data security while achieving efficient collaboration.
@@
As network technology continues to evolve, so do network attack methods, becoming increasingly sophisticated. conventional security mechanisms are often inadequate to counter these emerging threats. As shown in Figure 19, secure multi-party computation (SMPC) [205] is an encryption protocol that allows multiple parties to jointly compute a function while keeping their input data private, . Each party obtains the correct computation result without knowing the other parties’ private inputs. SMPC has characteristics such as decentralization, input data security, and accurate computation results. It enables AI to handle complex network environments, protect user privacy and data security, support emerging application scenarios, and promote data sharing and utilization.
Figure 19: Secure Multi-Party Computation [206]
@@
The main supporting technologies of SMPC include Garbled Circuits [207], Oblivious Transfer [208], Secret Sharing [209], and Homomorphic Encryption [210].
@@
(1) Garbled Circuits This technique involves compiling computational logic into a circuit and encrypting each gate (such as AND, XOR, etc.) within the circuit. Participants can complete the computation by interacting with the encrypted information without knowing the specific logic of the circuit.
@@
(2) Oblivious Transfer Oblivious Transfer is a protocol that allows one party (e.g., the sender) to send one of many pieces of information to another party (the receiver), where the receiver can choose only one piece of information. The sender does not know which piece of information the receiver has chosen.
@@
(3) Secret Sharing This method involves splitting a secret into multiple shares and distributing them to different participants. The original secret can only be reconstructed when a sufficient number of shares are combined.
@@
(4) Homomorphic Encryption Homomorphic Encryption is a special encryption method that allows specific computations (such as addition and multiplication) to be performed on encrypted data. The result of these computations, when decrypted, corresponds to the result of the same computations performed on the original data.
@@
SMPC, based on the aforementioned technologies, is widely applied in areas such as privacy-preserving ML. Recent research has leveraged SMPC to achieve secure model aggregation in FL, which involves training ML models on numerous decentralized devices while keeping the training data localized. This collaborative approach allows multiple parties to jointly train models without exposing their training data. For example, to counter reverse attacks on data, a privacy-preserving data aggregation mechanism based on secret sharing techniques in SMPC can efficiently protect FL against such attacks [211]. To address the issues of high communication overhead and poor scalability in conventional SMPC, a two-phase SMPC-based FL framework has been proposed [212]. This framework enables multiple clients to collaboratively train AI models while protecting their data privacy. To mitigate the high communication costs and poor scalability associated with aggregating global models at only some nodes, a hierarchical FL architecture using SMPC has been developed. This architecture improves communication efficiency and scalability without compromising privacy [213].
@@
In the future mobile network, data is often characterized by high volume, high frequency, and asynchronous communication, presenting new challenges for the implementation of FL. A novel hybrid FL architecture has been proposed, which combines FL with Trusted Execution Environments, SMPC and Beidou satellites. This architecture achieves secure key distribution, encryption, and decryption, and provides verification mechanisms for each participant to ensure the security of local data [214]. SMPC can also be combined with blockchain technology to track and mitigate malicious behavior, achieving a secure and trustworthy collaborative edge learning framework [215].
SMPC can protect data privacy while enabling collaborative computation among multiple parties, making it of significant and widespread application in the 6G network. SMPC offers an effective solution to the contradiction between data privacy protection and collaborative computing. It plays a crucial role in scenarios such as privacy-preserving FL, distributed resource allocation and optimization, secure data fusion and analysis, distributed identity verification and key management, and anti-fraud billing systems. By leveraging encrypted computation, SMPC ensures the secure processing of sensitive information. In FL, SMPC enables multiple devices or nodes to train models without sharing raw data, widely applied in scenarios like smart healthcare and intelligent cities. In resource allocation and optimization, SMPC protects sensitive data from each party, facilitating efficient collaboration in tasks such as spectrum resource allocation and edge computing load balancing. In data fusion and intelligent analysis, SMPC ensures the privacy of cross-domain data statistics and modeling, such as user behavior analysis and network traffic prediction. In identity verification and key management, SMPC mitigates the risk of single-point failures in centralized schemes, supporting secure access for IoT devices and edge computing nodes. In dynamic billing systems, SMPC guarantees the privacy and fairness of user billing calculations, applied in international roaming and dynamic pricing models.
Furthermore, SMPC also plays a role in content distribution optimization, network security protection, and intelligent supply chain management. By protecting data privacy, it enables cross-domain collaboration, such as optimizing content transmission paths, collaborative threat detection, and supply chain inventory management. Combined with technologies like homomorphic encryption and secret sharing, SMPC can address the privacy challenges of data sharing in the 6G network while optimizing protocol performance to meet the demands of low latency and high bandwidth. This not only enhances the privacy and security of the network but also drives the intelligent and distributed collaborative development of the 6G network.
@@ 3.2.5 Semantic Communication
Semantic communication, commonly relying on DNNs, is considered as a promising technology in 6G. Comparing with the traditional syntactic communication that focuses on the accurate transmission of data in bits, semantic communication only transmits the semantics of the message rather than the whole raw data. In this way, the amounts of data to be transmitted are largely reduced. Meanwhile, it can understand the data context, preserving essential semantic relationships when transmitting information, thereby aiding downstream AI systems in making more intelligent and accurate decisions. These characteristics are to meet the demands of some new emerging data-hungry applications in 6G, e.g., holographic communication, XR, show in Figure 20.
Figure 20: Some semantic communication enabled data-hungry applications in 6G
Figure 21 provides a basic semantic communication model, encompassing three essential components: the semantic encoder, the channel, and the semantic decoder. At the transmitter, semantic encoder extracts and encodes SI of the source data. Specifically, this process involves extracting the SI and compressing or removing the irrelevant information. To achieve this, the raw data is encoded by a neural network, donated as ${\mathcal{F}}_{\theta}$, which outputs the semantics that retains the critical meaning while discarding non-essential details. The compact semantics are then transmitted over a noisy physical channel, such as a wireless channel. As such, noises or interferences are introduced, thereby resulting in the semantic distortion. At the receiver, the semantic decoder decodes the received data. Specifically, this process involves “understanding” and inferring the semantic content sent by the transmitter. To accomplish this, the semantic decoder, donated as ${\mathcal{G}}_{\phi}$, processes the received data with the goal of mapping it back to the source data. This ensures that the reconstructed meaning accurately reflects the source message, even in the presence of potential transmission distortion.
In addition, 6G networks are full of multi-modal data from heterogeneous sources, such as text, image, audio, point cloud and etc. Despite the heterogenous data structure, the semantics behind the multi-modal data may tightly correlated under some conditions. As such, the semantic correlation between the data in different modalities support the data redundancy reduction further.
Figure 21: A basic semantic communication model
@@
The authors of [216] provided a pioneering work in SI theory, establishing the mathematical framework for SI theory, which lays the foundation for semantic communication. The concept of modulation division multiple access technology was proposed in [217], which efficiently distinguishes users based on semantic features from the information dimension, thereby improving resource utilization in multi-user semantic communication systems, which contributes to multiuser semantic communication systems. Semantic extraction and recovery, also called semantic encoding and decoding, are the core components of semantic transceivers. Four dominant advanced semantic extraction techniques [218]: DL-based semantic extraction, RL-based semantic extraction, KB-assisted semantic extraction, and native semantic extraction. Moreover, most semantic communication studies concentrate on text transmission, audio transmission and recognition, image transmission and recognition, and video transmission and recognition from the perspective of data modality and task classification.
@@
Text Transmission: The authors of [219] developed a DL-based text semantic communication system named DeepSC, which performs joint semantic-channel coding to achieve superior performance gains compared to existing technologies, particularly at low SNRs. Recognizing that most semantic metrics are non-differentiable, a RL-based optimization paradigm for large-scale and complex text semantic transmission was introduced in [220]. This technique uses self-critical random iterative updates to train decoupled semantic transceivers, addressing the non-differentiable semantic channel optimization issue. The authors of [221] proposed a text-oriented semantic communication technique named semantic coding Reed Solomon hybrid automatic repeat request (HARQ) by combining HARQ and Reed-Solomon channel coding with semantic coding to tackle inefficiency and inflexibility with varying sentence lengths.
@@
Audio Transmission and Recognition: A DL-based audio semantic communication system named DeepSC-S was developed, as noted in [222, 223]. The audio semantic communication scheme employs a joint semantic encoder/decoder and channel encoder/decoder to mitigate the semantic distortion caused by the noises and interference from the wireless channels. Additionally, the authors of [224] explored FL-based audio semantic communication, developing a wav2vec-based autoencoder composed of CNNs. This system effectively encodes, transmits, and decodes audio SI, reducing communication overhead.
@@
Image Transmission and Recognition: The authors of [225] investigated a DeepJSCC image semantic communication architecture, where the encoder and decoder functions are parameterized by CNNs and jointly trained on the same dataset to minimize the mean squared error of the reconstructed images. Joint training enables DeepJSCC to avoid the cliff effect and promptly adapts to the varying SNR.
@@
Video Transmission and Recognition: The authors of [226] developed a variable-length DeepJSCC system that utilizes nonlinear transformations and conditional coding architectures to adaptively extract the semantic features of video frames. This system outperforms the traditional wireless video coding transmission on recognition accuracy.
@@
Semantic communication has unique advantages in semantic understanding, driving advancements in applications such as the tactile internet, holographic communication, XR, and human-machine symbiotic networks, fostering an increasingly intelligent and efficient society.